# A characterization of the local structure of two-dimensional sets with positive reach

### v1

@@ 6. An alternative proof of Theorem L
In this section we present, using Theorem 4.1, an alternative proof of Lytchak’s Theorem L. Although the basic ideas of our approach come from [18], we present our proof since it is more elementary, as it does not need results from the theory of length spaces. Moreover, our proof of Proposition 6.1 (which is similar to the proof from [18, Subsection 5.2]) uses Lemma 2.23 (easily inferred from [8]) whereas the proof from [18] uses a (closely related) result on $O_{\delta}$ sets” from paper [24] which is accessible in Russian only.
@@
Denote $W:=e_{d}^{\perp}$. Applying [23, Lemma 3.5] and the definition of tangent vectors, we find that there exist $\tau>0$ and $L>0$ such that for any $w\in W\cap B(0,\tau)$,
@@
Consider now the case $1\leq k\leq d-1$. Then by Theorem 4.1 we can choose $\delta_{1}>0$ and $k$-dimensional $C^{1,1}$ surface $\Gamma$ such that $T_{k}(A)\cap B(a,\delta_{1})\subset\Gamma$. The assumptions together with Lemma 3.3 (i), Lemma 2.11 (iv) and Lemma 2.18 imply that there exists $0<\delta_{2}<\delta_{1}$ such that

### v2

@@ 6. An alternative proof of Theorem L
In this section we present, using Theorem 4.1, an alternative proof of Lytchak’s Theorem L. Although the basic ideas of our approach come from [17], we present our proof since it is more elementary, as it does not need results from the theory of length spaces. Moreover, our proof of Proposition 6.1 (which is similar to the proof from [17, Subsection 5.2]) uses Lemma 2.24 (easily inferred from [8]) whereas the proof from [17] uses a (closely related) result on $O_{\delta}$ sets” from paper [23] which is accessible in Russian only.
@@
Denote $W:=e_{d}^{\perp}$. Applying [22, Lemma 3.5] and the definition of tangent vectors, we find that there exist $\tau>0$ and $L>0$ such that for any $w\in W\cap B(0,\tau)$,
@@
Consider now the case $1\leq k\leq d-1$. Then by Theorem 4.1 (using also Definition 2.5) we can choose $\delta_{1}>0$ and $k$-dimensional $C^{1,1}$ g-surface $\Gamma$ such that $T_{k}(A)\cap B(a,\delta_{1})\subset\Gamma$. The assumptions together with Lemma 3.3 (i), Lemma 2.12 (iv) and Lemma 2.19 imply that there exists $0<\delta_{2}<\delta_{1}$ such that
@@
Acknowledgement
The authors are grateful to Alexander Lytchak for helpful comments on an earlier version of the manusript.